OpenStudy (anonymous):

How to solve cosx = sin(2x) on [0,2pi]

OpenStudy (anonymous):

so what is the question?

OpenStudy (anonymous):

It's part of a calculus question: Find the relative extrema of f(x) = 2sin(x)+cos(2x) on [0,2pi] using the second derivative test.

OpenStudy (anonymous):

cosx=2sinxcosx =>cosx(1-2sinx)=0 =>cosx=0 or 1/2=sinx so ans will be pi/2 ,3pi/2 ,5pi/6,pi/6

OpenStudy (anonymous):

I have derived the function to find the critical values: First derivative: F(x) = cos(x)-2sin(2x) Set to 0 and solve.... And we're back to my original question

OpenStudy (anonymous):

2cos(x)-2sin(2x)**

OpenStudy (anonymous):

Ahhh double angle formula! Thanks

OpenStudy (anonymous):

How did you get cos(1-2sinx)=0 though??

OpenStudy (campbell_st):

sin(2x) = 2sin(x)cos(x) so you will have cos(x) = 2cos(x)sin(x) then rewrite the equation equal to 0 and solve by fractoring